On Tuesday (Jan 12) the Economist magazine said that the exchange rate between the Canadian Dollar and the Euro was 1.55. On the same day, the Times newspaper said it was 0.65. Which got it wrong?

1
Expert's answer
2021-12-14T05:21:24-0500

The exchange rate ratio = the original currency ÷ new currency


Ratios 1.55 and 0.65 are related in terms of the exchange rate ratio used.

Ratio 1.55 is based on Canadian dollar/Euro, while ratio 0.65 is based on Euro/Canadian dollar


Canadian dollar/Euro = 155 ÷ 100 = 1.55


Euro/Canadian dollar = 100 ÷ 155 = 0.65


Therefore,the accurate figure is the one of 1.55 provided by the Economist magazine, indicating the Canadian dollar to Euro exchange rate.